OpenStudy (anonymous):

Calculate: cos(2arctan(1/2)) [Without the use of a calculator. :-) ]

jhonyy9 (jhonyy9):

like a first step write arctan = arcsin / arccos

OpenStudy (anonymous):

Does that relation hold?

jhonyy9 (jhonyy9):

hope you know that tan x = sin x / cos x and cot x = cos x / sin x yes ?

OpenStudy (anonymous):

Indeed

OpenStudy (amistre64):

arctan(1/2) = a; tan(a) = 1/2 = y/x, z=sqrt5 cos(2a) = cos^2(a) - sin^2(a) cos(a) = x/z, sin(a) = y/z

OpenStudy (anonymous):

Where do you derive z=sqrt5 from? D:

OpenStudy (amistre64):

but i cheated and used my fingers to "calculate" the z .... i got no idea how you would do it otherwise

OpenStudy (amistre64):

trig function are right triangles ...

OpenStudy (amistre64):

since x and y are defined by the tan; that leaves the pythag for the missing side

OpenStudy (amistre64):

2^2+1^2 = 5 , right?

OpenStudy (anonymous):

I'm quite certain you're supposed to do it through trigonometric relations etc; they're hinting with the application of the relations for cos(2x)

OpenStudy (amistre64):

i did do it thru trig relations ... just in a more basic trig relation process :)

OpenStudy (anonymous):

I don't follow. :(

OpenStudy (dls):

@Noliec The other way round, \[\Huge \cos2x=\frac{1-\tan^2x}{1+\tan^2x}\]

OpenStudy (dls):

substitute the values :)

OpenStudy (anonymous):

How do you derive this? :O

OpenStudy (amistre64):

not real sure how we could develop the cos(2x) = cos^2-sin^2 identity without some more fundamental processes

OpenStudy (anonymous):

Oh, I see what you did DLS! :D

OpenStudy (anonymous):

Thanks both of you! :) @DLS @amistre64

OpenStudy (dls):

easy to derive :) but you don't really need it just keep it in mind

OpenStudy (anonymous):

Really beautiful problem.